NI TCS bench works

Go down

NI TCS bench works

Post  Alessio on Mon Jan 25, 2010 5:07 pm

(Chiummo, Day, Di Paolo Emilio, Fafone, Morgia, Rocchi, Sperandio)

This is a summary of the works on the NI TCS bench of the last week.
First we tested the new CO2 lasers from Access Laser Co. We measured the output power: 22W for laser #1 and 28W for laser #2.
We placed the IR camera after the Brewster attenuator and measured the beam profiles (figure 1 for laser #1 and figure 2 for laser #2).
We also measured the open loop noise spectra of both lasers (figure 3): laser #2 seems a bit more noisy than #1, but this could depend on the particular laser temperature at the moment of the measurement. For both lasers a temperature scan is necessary to find a good working point.
We started by installing laser #2.
We re-aligned the whole bench, in particular we have placed back the lenses in their original position, since they were moved (NI CO2 laser changed).
At this point we noticed a strange behavior of the fixed polarizer that follows the half wave plate: rotating the polarizer by a very small angle (circa 1°) changes the shape of the transmitted beam (figures 4 and 5). The beam can become very elliptical, thus changing the power distribution and shape of the CO2 ring. The origin of this effect is still unclear and under investigation.
On friday, in order to find stable working points for the new NI CO2 laser, a temperature scan was launched (New NI laser temperature scan).









avatar
Alessio

Posts : 127
Join date : 2008-11-07
Age : 46
Location : Rome

View user profile

Back to top Go down

Back to top

- Similar topics

 
Permissions in this forum:
You cannot reply to topics in this forum